OpenStudy (anonymous):

what is 1/2 to the power of 100 - 89 will give medal

OpenStudy (alexandervonhumboldt2):

1/2^11=1/(2)^11=1/2048

OpenStudy (alexandervonhumboldt2):

1/2048

OpenStudy (anonymous):

thanks